OpenStudy (anonymous):

what is the square root of 5 over 11 minus the square root of 11 over 5

OpenStudy (anonymous):



OpenStudy (radar):

\[\sqrt{5}/\sqrt{11}-\sqrt{11}/\sqrt{5}\]\[\sqrt{5}\sqrt{5}/(\sqrt{55})-(\sqrt{11}\sqrt{11})/\sqrt{55}\]\[5/\sqrt{55} - 11/\sqrt{55}=-6/\sqrt{55}\]\[(-6/\sqrt{55})(\sqrt{55}/\sqrt{55}=-6\sqrt{55}/55=-.11\sqrt{55}=-.809\]